# A JWST Survey of the Supernova Remnant Cassiopeia A

### v1

@@ 4.1 Dust Emission
Cas A’s young age and proximity make it an ideal target for an in-depth investigation of SN dust formation and destruction processes. Herschel observations were able to probe the cold dust component in the inner unshocked ejecta for the first time (Barlow et al., 2010; De Looze et al., 2017; Priestley et al., 2019) and confirmed the high mass of newly formed dust ( M), but at a limited spatial resolution of . Similarly, the spatial resolution achieved with Spitzer IRS spectroscopy (2.5 to 6) was insufficient to resolve individual ejecta knots, which made it impossible to constrain the chemical pathways that lead to the formation of specific warm dust species in Cas A. Spitzer IRS did, however, identify the formation of many different dust populations around chemically distinct regions in Cas A (Rho et al., 2009; Arendt et al., 2014). JWST can associate spatially resolved warm dust knots with different chemical compositions (Ar, Ne, S, O) and has the sensitivity to detect distinct dust emission features in those individual knots. The detailed analysis of the MRS dust emission and decomposition of the various dust species that condensed in the ejecta of Cas A will be presented in a future publication. Here, we summarize the general properties of the dust emission observed with MRS and compare them to previous studies with Spitzer.
The MIRI/MRS spectra in Figure 2 all show a continuum from warm dust grains and various dust emission features. Position P1, which samples reverse-shocked ejecta with particularly bright argon emission, shows prominent broad dust features at 9, 12, and 21 µm. Arendt et al. (2014) attributed the 9 and 21 µm features to grains characterized by a relatively low magnesium-to-silicon ratio, while Rho et al. (2018) attributed the features to SiO grains. Arendt et al. (2014) argued that the 12 feature could arise from ellipsoidal SiO or SiC grains with a continuous distribution of axial ratios, and that the dust with this spectral profile is likely a mixture of silica and Mg silicates with Mg/Si 5. The MRS spectra show that this 12 µm feature is likely present in all four positions in Cas A, while the 9 and 21 µm features are only apparent in the argon-rich ejecta. The narrow dust feature at 11.3 that is likely produced by SiC grains (e.g. Jiang et al., 2005) is also present in all four spectra shown in Figure 2. We note that the same complex of dust emission features (9, 11.3, 12, and 21 ) is present in the spectrum of newly condensed SN dust in SNR G54.1+0.3 (Temim et al., 2010), which may be evidence that these features all arise from a common grain species or that the multiple grain species that produce the set of features are often found in the same environment.
@@ 4.2 Interior unshocked ejecta
The relative yields and distributions of elements, including radioactive nuclei freshly synthesized during the SN such as Ni and Ti (and their decay products) that originate from the innermost regions of the star where the explosion initiated, provide valuable constraints for these simulations (Grefenstette et al., 2014; Wongwathanarat et al., 2017; Grefenstette et al., 2017; Fryer et al., 2023; Sieverding et al., 2023). Hundreds of years after the SN, these nuclei continue to reflect the physical processes dominating the physics of SN engines and probe the degree of explosion asymmetry. Particularly important is the mass and distribution of Fe. The X-ray-bright reverse-shocked Fe-rich ejecta of Cas A (which trace the original Ni distribution) has velocities around and above 4000 km s (Hughes et al., 2000; DeLaney et al., 2010) and is associated with a mass of M (Hwang & Laming, 2012), which is a fair fraction of the total amount of Fe expected to have been ejected in a neutrino-driven explosion of about erg as estimated for Cas A (Orlando et al., 2016). For this reason, Cas A has been used as a clear example of how the ejecta asymmetries connected to the neutrino-driven mechanism and the nature of the progenitor as a H-stripped star can strongly influence the final distribution of ejecta by seeding Rayleigh-Taylor instabilities that in turn encourage radial mixing (Wongwathanarat et al., 2015, 2017).
@@
Presently, uncertainties in theoretical models (due to both the unknown exact mass and structure of the progenitor and uncertain degrees of freedom in the neutrino physics of supernova calculations) leave it unclear how much unshocked Fe is expected and observations have been unable to derive how much is actually present in Cas A (e.g., Young et al., 2006; Eriksen et al., 2009). Multiple lines of evidence support the notion that some Fe must still reside in the center of Cas A (see, e.g., Hwang & Laming 2012; Milisavljevic & Fesen 2015; Grefenstette et al. 2017). However, how much and how it is distributed are critical unknowns. By measuring the iron abundance, we obtain a better understanding of the nature of the central engine. Among the best efforts to date to investigate unshocked Fe was a deep 1.64 m image of Cas A by Koo et al. (2018), which likely traced interior [Si I] 1.645 m line emission as opposed to [Fe ii] 1.644 m line, and Laming & Temim (2020) that used Spitzer IRS spectra of Cas A. Laming & Temim (2020) estimated the total unshocked ejecta mass to be 0.47 M, which was broadly consistent with previous estimates (Hwang & Laming, 2012; DeLaney et al., 2014; Arias et al., 2018), and with mass fractions of 30% O, 60% Si, a few percent S, and traces of Ne and Ar. No Fe was confidently detected in the Spitzer data, but possible features at [Fe VIII] 5.446 and [Fe V] 20.85 m along with other considerations led to an upper limit Fe mass of M (Laming & Temim, 2020).
@@
We examined MIRI/MRS spectra of P2 and P4 closely for possible emission features around the [Fe VIII] 5.446 and [Fe V] 20.85 m emission lines. The only possible candidate is faint emission peaked around 5.39 m observed only in the P4 region potentially associated with blue-shifted [Fe VIII] 5.446 m. A second weaker emission peak is also observed at 5.53 m. The distribution of emission around these candidate lines is unlike other ejecta lines that are narrower and better resemble what would be expected from PAHs, though none are documented at these wavelengths. Nonetheless, if this emission originates from truly diffuse, extended gas, then the observed distribution is viable. Following assumptions in Laming & Temim (2020) that emission lines come from density regions where their emission is maximized, we estimate that the putative [Fe viii] emission in the P4 field of view (approximately ergs cm s sr) represents M of Fe. If replicated similarly across the remnant, this would imply M of unshocked Fe material. However, since the candidate emission feature is absent from P2 spectra, clearly the Fe is not distributed uniformly. Future work can investigate the reality of this candidate detection and carefully utilize all available images and spectra to develop the best constraint possible of the total Fe yield from the supernova.
@@
We note that portions of our unshocked ejecta map may be contaminated by cold dust embedded in the interior ejecta, as well as emission from filaments of shocked ejecta. However, many arguments strongly support the view that the dominant emission is indeed from unshocked ejecta. For example, the bright “Central arc” is consistent with interior structure already mapped in [S iii] 9069, 9531 by Milisavljevic & Fesen (2015) having a velocity distribution of to km s. Likewise, the distribution of emission in our unshocked ejecta map strongly correlates with features interior to the main shell only seen in the [S iii]-sensitive F1000W image.
The centrally-located P2 and P4 MIRI/MRS spectra provide a complementary probe of the unshocked ejecta through the center of the remnant. We isolated [O iv] emission in each spaxal by exploring the data cube with the STScI-developed JDAViz tool (JDADF Developers et al., 2023) to find appropriate wavelength regions to model the continuum. A 3rd order polynomial was used for the continuum fit, which was then subtracted from the data cube. The cube was cropped by pixels on each side to eliminate artifacts from the edge of the detector. Coordinates in right ascension and declination were converted into velocity using the scaling factor of 0.022 per km s (Milisavljevic & Fesen, 2013), with respect to the center of expansion (Thorstensen et al., 2001). The data were linearly interpolated and resampled with a resolution of 22 km s in all three dimensions, which is comparable to the original spatial resolution.
The continuum-subtracted, cropped, interpolated [O iv] emission from P2 and P4 as a function of velocity is shown in the bottom of Figure 4 from two viewing angles. Previous observations by Spitzer using the IRS presented by Isensee et al. (2010) with spatial resolution of and spectral resolution of 0.05 m indicated sheet-like structures and filaments, including holes within the sheets at the scale of 30″. Our data reveal an even more intricate distribution, with filaments seen at the scale of 1″. The velocities range over 3000 km s, indicative of ejecta velocities and not CSM. We interpret these MIRI/MRS as further confirmation that the interior network of filaments seen in the top panel in Figure 4 traces unshocked ejecta accurately.
@@ 5 Light echoes
Echoes of various brightness are found across the field of our JWST observations. Figure 5 highlights two prominent regions of echoes. The region labeled E2 shows typical bright echoes, which appear as bright clumps embedded in fainter more filamentary structures. The additional JWST observations discussed in section 2 were performed 109 days following the initial observations, providing an unexpected opportunity to search for variations of light echo features surrounding portions of Cas A. The lower right panel in Figure 5 shows the dramatic changes seen in the MIRI F770W images between these two epochs.
@@
The incredible structure in this echo, showing many small isolated features and concentric arclets, along with its overall well defined outer edge, indicates a very complex ISM cloud surface. The thinnest strands remain unresolved in width even in our NIRCam F162M image. To our knowledge, this is the best view yet of ISM structure at the smallest of scales possible: 0.1 translates to A.U., or light-days, which is consistent with the echoes changing completely on time scales of months (Dwek & Arendt, 2008), and implies that changes on time scales of hours are possible. The F162M was observed in parallel with both the F356W and F444W observations, which occurred 19.3 hours apart in the initial mapping. Comparison of the F162M images of the large IR echo at these epochs does not reveal any very short term changes in the echo brightness. This suggests that there were no major changes in the illuminating SN emission on this time scale and/or that the echoing structures are lt-hr A.U. in size. However, F162M is not the band that is most sensitive to the echo emission.
@@ 6 Multi-wavelength comparisons
Cas A is one of the brightest radio sources on the sky (Reber, 1944), having a total flux density of Jy in the 1-2 GHz band (Baars et al., 1977). A relatively circular region with radius is associated with the reverse shock propagating into the expanding ejecta. Weaker radio emission associated with synchrotron radiation (Ginzburg & Syrovatskii, 1965) is seen at some azimuths extending out to the location of the main forward shock, traveling at 4200 – 5200 km s (DeLaney & Rudnick, 2003; Patnaude & Fesen, 2009).
@@
In Figure 6 (lower left panel), we show a composite image using 0.3-10 keV Chandra broadband observations (Hwang & Laming, 2012), a NuSTAR Ti map (Grefenstette et al., 2014), and our MIRI F1280W image. A gap in the Ti emission that runs north-south is approximately where the brightest ridge of the Green Monster is located. Because we favor the interpretation that the bulk emission of the Green Monster is associated with CSM on the front side of the remnant (towards us), the anti-correlation with Ti (which is redshifted overall; Grefenstette et al. 2017) should be coincidental. Absorption of Ti emission ( keV) from the Green Monster seems unlikely, since a CSM column density of cm would be needed to reduce the flux by a factor of 10. Nonetheless, clear correlations in emission are seen between Chandra and our JWST data, suggesting a shared origin. De Looze et al. (2024) and Vink et al. (2024) investigate this X-ray–IR relationship in detail and attribute it to interaction between the forward shock and relatively dense CSM.
Another composite image is shown in Figure 6 (lower right) that compares our unshocked ejecta map (section 4.2; Figure 4) with a Chandra Fe–K map (energy keV; tracing where Ni was produced), and the NuSTAR Ti map. As noted earlier, the SE filaments and Western branch of the unshocked ejecta connect with Fe-rich locations. On the other hand, no strong relationship is seen between [O iv] and Ti, except that Ti emission is located where [O iv] is weak. Together, these comparisons support the view that the fine structures in the interior volume are most likely rooted in turbulent mixing of cool, low-entropy matter from the progenitor’s oxygen layer with hot, neutrino and radioactively heated high-entropy matter.
@@ 7 Neutron Star
Non-detections outside the X-ray range are usually attributed to a lack of pulsar activity because of the low magnetic field or to high extinction (Pavlov et al., 2004; Halpern & Gotthelf, 2010; De Luca, 2017). Our NIRCam images offer the possibility to peer deeply enough through high extinction at wavelengths other than X-rays to test hypotheses about the CCO’s nature. To this end, we conducted forced photometry at the coordinates of the CCO ((J2000) = 232327943, (J2000) = +58425; Fesen et al. 2006b), on the three NIRCam filter images using the space_phot222See https://app.soos.io/research/packages/Python/-/space-phot. package. The photometry was measured using point-spread function (PSF) photometry on the level 2 NIRCam “CAL” images, which preserves the PSF structure compared to level 3 resampled images. The NIRCam pixel area map for the corresponding SW and LW modules must also be applied to each exposure to correct for pixel area variations across the images. Here we use the PSF models provided by WebbPSF (version 1.2.1)333https://www.stsci.edu/jwst/science-planning/proposal-planning-toolbox/psf-simulation-tool to represent the PSF, which have been updated to better match the observed PSF in each filter, and take into account temporal and spatial variation across the detector. We remove a constant background in the fitting region, which is small enough that a constant background is sufficient. The background is estimated as the median value in an annulus centered at the position with an inner radius of 5 pixels and outer radius 7 pixels. space_phot simultaneously constrains the common flux in all CALs at the forced photometry position, within a pixel square. The final measured flux is the integral of each full fitted PSF model, which includes a correction to the infinite aperture flux. These total fluxes, which are in units of MJy/sr, are converted to AB magnitudes using the native pixel scale of each image.
@@
Firstly, we can constrain the area of any circular disk with a blackbody temperature similar to the one inferred from the Spitzer detection of the magnetar 4U 0142+61 (Wang et al., 2006). Such a disk with temperature K must have an area smaller than cm (i.e., ), where is the disk inclination, and is the outer radius of the disk.

### v2

@@ 4.1 Dust Emission
Cas A’s young age and proximity make it an ideal target for an in-depth investigation of SN dust formation and destruction processes. Herschel observations were able to probe the cold dust component in the inner unshocked ejecta for the first time (Barlow et al., 2010; De Looze et al., 2017; Priestley et al., 2019) and confirmed the high mass of newly formed dust ( M⊙), but at a limited spatial resolution of . Similarly, the spatial resolution achieved with Spitzer IRS spectroscopy (2.5′′ to 6′′) was insufficient to resolve individual ejecta knots, which made it impossible to constrain the chemical pathways that lead to the formation of specific warm dust species in Cas A. Spitzer IRS did, however, identify the formation of many different dust populations around chemically distinct regions in Cas A (Rho et al., 2009; Arendt et al., 2014). JWST can associate spatially resolved warm dust knots with different chemical compositions (Ar, Ne, S, O) and has the sensitivity to detect distinct dust emission features in those individual knots. The detailed analysis of the MRS dust emission and decomposition of the various dust species that condensed in the ejecta of Cas A will be presented in a future publication. Here, we summarize the general properties of the dust emission observed with MRS and compare them to previous studies with Spitzer.
The MIRI/MRS spectra in Figure 2 all show a continuum from warm dust grains and various dust emission features. Position P1, which samples reverse-shocked ejecta with particularly bright argon emission, shows prominent broad dust features at 9, 12, and 21 µm. Arendt et al. (2014) attributed the 9 and 21 µm features to grains characterized by a relatively low magnesium-to-silicon ratio, while Rho et al. (2018) attributed the features to SiO2 grains. Arendt et al. (2014) argued that the 12 feature could arise from ellipsoidal SiO2 or SiC grains with a continuous distribution of axial ratios, and that the dust with this spectral profile is likely a mixture of silica and Mg silicates with Mg/Si 5. The MRS spectra show that this 12 µm feature is likely present in all four positions in Cas A, while the 9 and 21 µm features are only apparent in the argon-rich ejecta. The narrow dust feature at 11.3 that is likely produced by SiC grains (e.g. Jiang et al., 2005) is also present in all four spectra shown in Figure 2. We note that the same complex of dust emission features (9, 11.3, 12, and 21 ) is present in the spectrum of newly condensed SN dust in SNR G54.1+0.3 (Temim et al., 2010), which may be evidence that these features all arise from a common grain species or that the multiple grain species that produce the set of features are often found in the same environment.
@@ 4.2 Interior unshocked ejecta
The relative yields and distributions of elements, including radioactive nuclei freshly synthesized during the SN such as 56Ni and 44Ti (and their decay products) that originate from the innermost regions of the star where the explosion initiated, provide valuable constraints for these simulations (Grefenstette et al., 2014; Wongwathanarat et al., 2017; Grefenstette et al., 2017; Fryer et al., 2023; Sieverding et al., 2023). Hundreds of years after the SN, these nuclei continue to reflect the physical processes dominating the physics of SN engines and probe the degree of explosion asymmetry. Particularly important is the mass and distribution of Fe. The X-ray-bright reverse-shocked Fe-rich ejecta of Cas A (which trace the original 56Ni distribution) has velocities around and above 4000 km s-1 (Hughes et al., 2000; DeLaney et al., 2010) and is associated with a mass of M⊙ (Hwang & Laming, 2012), which is a fair fraction of the total amount of Fe expected to have been ejected in a neutrino-driven explosion of about erg as estimated for Cas A (Orlando et al., 2016). For this reason, Cas A has been used as a clear example of how the ejecta asymmetries connected to the neutrino-driven mechanism and the nature of the progenitor as a H-stripped star can strongly influence the final distribution of ejecta by seeding Rayleigh-Taylor instabilities that in turn encourage radial mixing (Wongwathanarat et al., 2015, 2017).
@@
Presently, uncertainties in theoretical models (due to both the unknown exact mass and structure of the progenitor and uncertain degrees of freedom in the neutrino physics of supernova calculations) leave it unclear how much unshocked Fe is expected and observations have been unable to derive how much is actually present in Cas A (e.g., Young et al., 2006; Eriksen et al., 2009). Multiple lines of evidence support the notion that some Fe must still reside in the center of Cas A (see, e.g., Hwang & Laming 2012; Milisavljevic & Fesen 2015; Grefenstette et al. 2017). However, how much and how it is distributed are critical unknowns. By measuring the iron abundance, we obtain a better understanding of the nature of the central engine. Among the best efforts to date to investigate unshocked Fe was a deep 1.64 m image of Cas A by Koo et al. (2018), which likely traced interior [Si I] 1.645 m line emission as opposed to [Fe II] 1.644 m line, and Laming & Temim (2020) that used Spitzer IRS spectra of Cas A. Laming & Temim (2020) estimated the total unshocked ejecta mass to be 0.47 M⊙, which was broadly consistent with previous estimates (Hwang & Laming, 2012; DeLaney et al., 2014; Arias et al., 2018), and with mass fractions of 30% O, 60% Si, a few percent S, and traces of Ne and Ar. No Fe was confidently detected in the Spitzer data, but possible features at [Fe VIII] 5.446 and [Fe V] 20.85 m along with other considerations led to an upper limit Fe mass of M⊙ (Laming & Temim, 2020).
@@
We examined MIRI/MRS spectra of P2 and P4 closely for possible emission features around the [Fe VIII] 5.446 and [Fe V] 20.85 m emission lines. The only possible candidate is faint emission peaked around 5.39 m observed only in the P4 region potentially associated with blue-shifted [Fe VIII] 5.446 m. A second weaker emission peak is also observed at 5.53 m. The distribution of emission around these candidate lines is unlike other ejecta lines that are narrower and better resemble what would be expected from PAHs, though none are documented at these wavelengths. Nonetheless, if this emission originates from truly diffuse, extended gas, then the observed distribution is viable. Following assumptions in Laming & Temim (2020) that emission lines come from density regions where their emission is maximized, we estimate that the putative [Fe VIII] emission in the P4 field of view (approximately ergs cm-2 s-1 sr-1) represents M⊙ of Fe. If replicated similarly across the remnant, this would imply M⊙ of unshocked Fe material. However, since the candidate emission feature is absent from P2 spectra, clearly the Fe is not distributed uniformly. Future work can investigate the reality of this candidate detection and carefully utilize all available images and spectra to develop the best constraint possible of the total Fe yield from the supernova.
@@
We note that portions of our unshocked ejecta map may be contaminated by cold dust embedded in the interior ejecta, as well as emission from filaments of shocked ejecta. However, many arguments strongly support the view that the dominant emission is indeed from unshocked ejecta. For example, the bright “Central arc” is consistent with interior structure already mapped in [S III] 9069, 9531 by Milisavljevic & Fesen (2015) having a velocity distribution of to km s-1. Likewise, the distribution of emission in our unshocked ejecta map strongly correlates with features interior to the main shell only seen in the [S IV]-sensitive F1000W image.
The centrally-located P2 and P4 MIRI/MRS spectra provide a complementary probe of the unshocked ejecta through the center of the remnant. We isolated [O IV] emission in each spaxal by exploring the data cube with the STScI-developed JDAViz tool (JDADF Developers et al., 2023) to find appropriate wavelength regions to model the continuum. A 3rd order polynomial was used for the continuum fit, which was then subtracted from the data cube. The cube was cropped by pixels on each side to eliminate artifacts from the edge of the detector. Coordinates in right ascension and declination were converted into velocity using the scaling factor of 0.022′′ per km s-1 (Milisavljevic & Fesen, 2013), with respect to the center of expansion (Thorstensen et al., 2001). The data were linearly interpolated and resampled with a resolution of 22 km s-1 in all three dimensions, which is comparable to the original spatial resolution.
The continuum-subtracted, cropped, interpolated [O IV] emission from P2 and P4 as a function of velocity is shown in the bottom of Figure 4 from two viewing angles. Previous observations by Spitzer using the IRS presented by Isensee et al. (2010) with spatial resolution of and spectral resolution of 0.05 m indicated sheet-like structures and filaments, including holes within the sheets at the scale of 30″. Our data, which localize ejecta velocities with uncertainty 160 km s-1 for any 3D location of the [O IV] emission and improve over Spitzer observations by approximately an order of magnitude, reveal an even more intricate distribution with filaments seen at the scale of 1″. The velocities range over 3000 km s-1, indicative of ejecta and not CSM. We interpret the MIRI/MRS spectra as further confirmation that the interior network of filaments seen in the top panel in Figure 4 traces unshocked ejecta accurately.
@@ 5 Light echoes
Echoes of various brightness are found across the field of our JWST observations. Figure 5 highlights two prominent regions of echoes. The region labeled E2 shows typical bright echoes, which appear as bright clumps embedded in fainter more filamentary structures. The additional JWST observations discussed in section 2 were performed 109 days following the initial observations, providing an opportunity to search for variations of light echo features surrounding portions of Cas A. The lower right panel in Figure 5 shows the dramatic changes seen in the MIRI F770W images between these two epochs. These can be confidently identified as light echoes since the proper motion of emission, if interpreted as the same clumps of material from one epoch to the next, would imply unrealistic velocities between 0.3c to 0.6c. The fastest moving ejecta in Cas A is traveling at 0.05c in the outermost portion of the northeast jet (Fesen & Milisavljevic, 2016).
@@
The incredible structure in this echo, showing many small isolated features and concentric arclets, along with its overall well defined outer edge, indicates a very complex ISM cloud surface. The thinnest strands remain unresolved in width even in our NIRCam F162M image. To our knowledge, this is the best view yet of ISM structure at the smallest of scales possible: 0.1′′ translates to A.U., or light-days, which is consistent with the echoes changing completely on time scales of months (Dwek & Arendt, 2008), and implies that changes on time scales of hours are possible. The F162M was observed in parallel with both the F356W and F444W observations, which occurred 19.3 hours apart in the initial mapping. Comparison of the F162M images of the large IR echo at these epochs does not reveal any very short term changes in the echo brightness. This suggests that there were no major changes in the illuminating SN emission on this time scale and/or that the echoing structures are lt-hr A.U. in size. However, F162M is not the band that is most sensitive to the echo emission.
@@ 6 Multi-wavelength comparisons
Cas A is one of the brightest radio sources on the sky (Reber, 1944), having a total flux density of Jy in the 1-2 GHz band (Baars et al., 1977). A relatively circular region with radius is associated with the reverse shock propagating into the expanding ejecta. Weaker radio emission associated with synchrotron radiation (Ginzburg & Syrovatskii, 1965) is seen at some azimuths extending out to the location of the main forward shock, traveling at 4200 – 5200 km s-1 (DeLaney & Rudnick, 2003; Patnaude & Fesen, 2009).
@@
In Figure 6 (lower left panel), we show a composite image using 0.3-10 keV Chandra broadband observations (Hwang & Laming, 2012), a NuSTAR 44Ti map (Grefenstette et al., 2014), and our MIRI F1280W image. A gap in the 44Ti emission that runs north-south is approximately where the brightest ridge of the Green Monster is located. Because we favor the interpretation that the bulk emission of the Green Monster is associated with CSM on the front side of the remnant (towards us), the anti-correlation with 44Ti (which is redshifted overall; Grefenstette et al. 2017) should be coincidental. Absorption of 44Ti emission ( keV) from the Green Monster seems unlikely, since a CSM column density of cm-2 would be needed to reduce the flux by a factor of 10. Nonetheless, clear correlations in emission are seen between Chandra and our JWST data, suggesting a shared origin. De Looze et al. (in preparation) and Vink et al. (2024) investigate this X-ray–IR relationship in detail and attribute it to interaction between the forward shock and relatively dense CSM.
Another composite image is shown in Figure 6 (lower right) that compares our unshocked ejecta map (section 4.2; Figure 4) with a Chandra Fe–K map (energy keV; tracing where 56Ni was produced), and the NuSTAR 44Ti map. As noted earlier, the SE filaments and Western branch of the unshocked ejecta connect with Fe-rich locations. On the other hand, no strong relationship is seen between [O IV] and 44Ti, except that 44Ti emission is located where [O IV] is weak. Together, these comparisons support the view that the fine structures in the interior volume are most likely rooted in turbulent mixing of cool, low-entropy matter from the progenitor’s oxygen layer with hot, neutrino and radioactively heated high-entropy matter.
@@ 7 Neutron Star
Non-detections outside the X-ray range are usually attributed to a lack of pulsar activity because of the low magnetic field or to high extinction (Pavlov et al., 2004; Halpern & Gotthelf, 2010; De Luca, 2017). Our NIRCam images offer the possibility to peer deeply enough through high extinction at wavelengths other than X-rays to test hypotheses about the CCO’s nature. To this end, we conducted forced photometry at the coordinates of the CCO ((J2000) = 23h23m27943, (J2000) = +58o425; Fesen et al. 2006b), on the three NIRCam filter images using the space_phot222See https://app.soos.io/research/packages/Python/-/space-phot. package. The photometry was measured using point-spread function (PSF) photometry on the level 2 NIRCam “CAL” images, which preserves the PSF structure compared to level 3 resampled images. The NIRCam pixel area map for the corresponding SW and LW modules must also be applied to each exposure to correct for pixel area variations across the images. Here we use the PSF models provided by WebbPSF (version 1.2.1)333https://www.stsci.edu/jwst/science-planning/proposal-planning-toolbox/psf-simulation-tool to represent the PSF, which have been updated to better match the observed PSF in each filter, and take into account temporal and spatial variation across the detector. We remove a constant background in the fitting region, which is small enough that a constant background is sufficient. The background is estimated as the median value in an annulus centered at the position with an inner radius of 5 pixels and outer radius 7 pixels. space_phot simultaneously constrains the common flux in all CALs at the forced photometry position, within a pixel square. The final measured flux is the integral of each full fitted PSF model, which includes a correction to the infinite aperture flux. These total fluxes, which are in units of MJy/sr, are converted to AB magnitudes using the native pixel scale of each image.
@@
Firstly, we can constrain the area of any circular disk with a blackbody temperature similar to the one inferred from the Spitzer detection of the magnetar 4U 0142+61 (Wang et al., 2006). Such a disk with temperature K must have an area smaller than cm2 (i.e., ), where is the disk inclination, and is the outer radius of the disk.